OpenStudy (alyssajobug):

Use f(t) = 18 - 2.33s for Black Dragon and f(t) = 20 - 3.06s for Wild Thing. After how many seconds would Black Dragon and Wild Thing have the same amount of fuel left in their tanks? Round your answer to the nearest hundredth of a second.

OpenStudy (texaschic101):

18 - 2.33s = 20 - 3.06s 18 - 20 = -3.06s + 2.33s -2 = - 0.73s -2/-0.73 = s 2.739 = s rounded to nearest hundredth = 2.74 seconds